# Octax: Accelerated CHIP-8 Arcade Environments for Reinforcement Learning in JAX

### v1

@@ 3.3 How to transform games into RL environments?
We address CHIP-8’s non-standardized scoring and termination by combining static ROM analysis and dynamic memory monitoring during gameplay, as detailed in Appendix C.
@@ 3.4 Which games does Octax support?
The games (Figure 1) vary across multiple dimensions of difficulty and cognitive demand. Temporal complexity ranges from immediate reactions to long-term planning requirements. Spatial complexity spans single-screen environments to multi-screen worlds requiring navigation. Reward structures include both dense scoring mechanisms and sparse achievement-based systems. This systematic variation enables controlled studies of algorithmic performance across different challenge types while maintaining a unified technical framework for fair comparison. A categorization of these games is provided in Table 1, with more detailed descriptions available in Appendix C.2.
@@ 4.3 How do LLMs assist environment creation?
Automated environment generation pipeline. Our pipeline consists of seven replicable steps for automatic CHIP-8 game generation. In Step 1, we construct a corpus of CHIP-8 tutorials, documentation, and programming examples, ensuring the LLM understands the architecture’s instruction set, memory layout, and common coding patterns. In Step 2, we embed this corpus into a prompt (detailed in Appendix E.1) that guides the LLM to produce syntactically correct CHIP-8 programs from high-level instructions. In Step 3, we provide a description of the game with desired mechanics, objectives, and constraints. In Step 4, the LLM generates the initial CHIP-8 code based on the provided description. In Step 5, an automated feedback loop between the LLM and a CHIP-8 compiler iteratively refines the code based on compilation errors until successful. In Step 6, Python wrapper functions for score_fn and terminated_fn are automatically generated, translating CHIP-8 registers into RL-compatible reward and termination signals. Finally, in Step 7, the game description is augmented to increase difficulty or introduce new challenges. Both the new description and the previously generated game are added to the LLM’s context before next iteration. Figure 7 summarizes the automated environment generation pipeline.
Target Shooter case study. We validated this pipeline using Claude Opus 4.1, known for its proficiency in programming, with the following description: "Target Shooter – Targets appear randomly on the screen, and the player moves a crosshair to shoot them. Score increases per hit, and the game ends after a fixed number of targets." The system successfully generated three progressive difficulty levels: static targets for basic aiming skills, time-limited targets introducing decision pressure, and moving targets with time constraints requiring predictive aiming. Each level maintains consistent register mappings for score and termination, simplifying Octax compatibility. Figure 6 shows how the LLM-generated environment visual appearance. All the code generated by the LLM is given in Appendix E.2,
@@ 5 Conclusion
Reproducibility Statement
We provide complete resources to ensure reproducibility of our results. The Octax source code, including all 21 game environment implementations, JAX-based CHIP-8 emulator, and training scripts, is available as supplementary material. Our experimental setup uses standard PPO hyperparameters detailed in Section 4.1, with hardware specifications and performance benchmarking configurations provided in Section 4.2. All training experiments use identical network architectures and hyperparameters across games, enabling direct replication of our learning curves in Figure 3. For the LLM-assisted environment generation pipeline in Section 4.3, we include the prompt templates and generated CHIP-8 assembly code in Appendix E. The modular design of Octax allows researchers to extend our game suite using the technical specifications in Section 3. The repository containing all source code, experiments, and data is available at: https://github.com/riiswa/octax.
Appendix A Use of Large Language Models
We used Large Language Models in three capacities during this research. First, Claude Opus 4.1 serves as a core research component in Section 4.3, where we demonstrate automated CHIP-8 game generation from high-level descriptions. This represents a novel research contribution, with all generated code validated through compilation and RL experiments. Second, we employed Claude Sonnet 4 for writing assistance, including text refinement, rephrasing technical concepts, and improving academic tone. Third, LLMs generated code documentation, docstrings, and tutorial content. All research ideas, experimental design, and scientific claims originate from the authors. We did not use LLMs for ideation, hypothesis formation, or result interpretation. We manually reviewed and validated all LLM-assisted content for accuracy and take full responsibility for all presented content.
Appendix B CHIP-8 Technical Specifications
B.1 Platform Overview
@@ B.4 Font System
Appendix C Game Environment Implementation Details
C.1 Score Detection Methodology
@@
C.2 Game List
C.2.1 Long-horizon Planning & Spatial Reasoning
@@
C.2.2 Timing, Prediction & Reactive Control
@@
C.2.3 Resource Management & Tactical Decisions
@@
C.2.4 Exploration & Continuous Navigation
@@
C.2.5 Simple Reaction & Timing
@@
Appendix D Hyperparameter Optimization Results
@@ D.4 Final Configuration
Appendix E LLM-Assisted Environment Generation
@@ E.1 Prompt Engineering
E.2 Generated Target Shooter Implementation
@@
E.2.1 Level 1: Static Targets
@@
E.2.2 Level 2: Time-Limited Targets
@@
E.2.3 Level 3: Moving Targets with Time Constraints
@@
E.2.4 Environment Integration and Wrapper Implementation

### v2

@@ 3.3 How to transform games into RL environments?
We address CHIP-8’s non-standardized scoring and termination by combining static ROM analysis and dynamic memory monitoring during gameplay, as detailed in Appendix B.
@@ 3.4 Which games does Octax support?
The games (Figure 1) vary across multiple dimensions of difficulty and cognitive demand. Temporal complexity ranges from immediate reactions to long-term planning requirements. Spatial complexity spans single-screen environments to multi-screen worlds requiring navigation. Reward structures include both dense scoring mechanisms and sparse achievement-based systems. This systematic variation enables controlled studies of algorithmic performance across different challenge types while maintaining a unified technical framework for fair comparison. A categorization of these games is provided in Table 1, with more detailed descriptions available in Appendix B.2.
@@ 4.3 How do LLMs assist environment creation?
Automated environment generation pipeline. Our pipeline consists of seven replicable steps for automatic CHIP-8 game generation. In Step 1, we construct a corpus of CHIP-8 tutorials, documentation, and programming examples, ensuring the LLM understands the architecture’s instruction set, memory layout, and common coding patterns. In Step 2, we embed this corpus into a prompt (detailed in Appendix D.1) that guides the LLM to produce syntactically correct CHIP-8 programs from high-level instructions. In Step 3, we provide a description of the game with desired mechanics, objectives, and constraints. In Step 4, the LLM generates the initial CHIP-8 code based on the provided description. In Step 5, an automated feedback loop between the LLM and a CHIP-8 compiler iteratively refines the code based on compilation errors until successful. In Step 6, Python wrapper functions for score_fn and terminated_fn are automatically generated, translating CHIP-8 registers into RL-compatible reward and termination signals. Finally, in Step 7, the game description is augmented to increase difficulty or introduce new challenges. Both the new description and the previously generated game are added to the LLM’s context before next iteration. Figure 7 summarizes the automated environment generation pipeline.
Target Shooter case study. We validated this pipeline using Claude Opus 4.1, known for its proficiency in programming, with the following description: "Target Shooter – Targets appear randomly on the screen, and the player moves a crosshair to shoot them. Score increases per hit, and the game ends after a fixed number of targets." The system successfully generated three progressive difficulty levels: static targets for basic aiming skills, time-limited targets introducing decision pressure, and moving targets with time constraints requiring predictive aiming. Each level maintains consistent register mappings for score and termination, simplifying Octax compatibility. Figure 6 shows how the LLM-generated environment visual appearance. All the code generated by the LLM is given in Appendix D.2,
@@ 5 Conclusion
Appendix A CHIP-8 Technical Specifications
A.1 Platform Overview
@@ A.4 Font System
Appendix B Game Environment Implementation Details
B.1 Score Detection Methodology
@@
B.2 Game List
B.2.1 Long-horizon Planning & Spatial Reasoning
@@
B.2.2 Timing, Prediction & Reactive Control
@@
B.2.3 Resource Management & Tactical Decisions
@@
B.2.4 Exploration & Continuous Navigation
@@
B.2.5 Simple Reaction & Timing
@@
Appendix C Hyperparameter Optimization Results
@@ C.4 Final Configuration
Appendix D LLM-Assisted Environment Generation
@@ D.1 Prompt Engineering
D.2 Generated Target Shooter Implementation
@@
D.2.1 Level 1: Static Targets
@@
D.2.2 Level 2: Time-Limited Targets
@@
D.2.3 Level 3: Moving Targets with Time Constraints
@@
D.2.4 Environment Integration and Wrapper Implementation